OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3 My options^

OpenStudy (anonymous):

@hartnn @jim_thompson5910 ?

hartnn (hartnn):

line that passes through the point (x1,y1) and has a slope of m is \(y-y_1=m(x-x_1)\) then,line that passes through the point (−1, 6) and has a slope of −3 will be ...?

OpenStudy (anonymous):

Oh, I get it now, haha. It's y − 6 = −3(x + 1) , right?

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

Woohoo! Thank you :)